the demand for ice cream is x=10+m/10p. the income of the consumer is $120 and the price of the ice cream is $3 . now suppose the price of the ice cream falls to $2 . find the income and substitution effect.
1
Expert's answer
2020-06-29T15:33:46-0400

"X=10+\frac{m}{10p}"

His demand for ice cream will be;

"10+\frac{120}{30}"

"=10+4=14 dollars"


If the ice cream decreases by 2 dollars then the new price will be;

"10+\frac{120}{20}"

"=10+6=16" dollars

The total change in demand will be "16-14=2" dollars

"\Delta""X"15"=X"1"(P"1",m)-X"1"(P"1",m)"

"\Delta""m=X"1 "\Delta" "P"1"=14(2-3)=" -"14" dollars

Therefore consumers income will be reduced to 14 dollars in order to hold his purchasing power.

"=m+\Delta" "m=120-14=106" dollars.

"\text{With an income of 6 he can still purchase 14 units of the ice cream at a lower 
price of 2 dollars}""\text{Consumers demand for the ice cream when he faces the price of 2 dollars and has an income of 106 dollars}" "=10+\frac{106}{20}" "=15.3" dollars.

The substitution effect will therefore be ;

"\Delta""X"51"=X"1"(2,106)-" "X"1"(3,120)"

"=15.3-14=1.3" dollars.This is the substitution effect.

The income effect is;

"\Delta""X"n1"=""X"1"(P"1",m)-" "X"1"(P"1",m)"

"\Delta""X"n1"=X"1"(2,120)-" "X"1"(2,106)"

"=16-15.3=0.7" dollars

Therefore the income effect is "0.7" "dollars"